OpenStudy (anonymous):

The refractive index of cooking oil is 1.47, and the refractive index of water is 1.33. A thick layer of cooking oil is floating on top of water in a beaker. What is the critical angle for light passing from the oil into the water? 25º, 43º , 49º , or 65º ?

OpenStudy (michele_laino):

we have to apply the law of Snell, so we can write: \[\Large {n_{OIL}}\sin {\theta _{OIL}} = {n_{WATER}} \cdot 1\]

OpenStudy (anonymous):

ok!

OpenStudy (michele_laino):

so: \[\Large \sin {\theta _{OIL}} = \frac{{{n_{WATER}}}}{{{n_{OIL}}}} = \frac{{1.33}}{{1.47}}\]

OpenStudy (anonymous):

0.90476 ?

OpenStudy (michele_laino):

and what is: arcsin(0.90476)=...?

OpenStudy (anonymous):

0.0157916973 ?

OpenStudy (michele_laino):

no, it is : 64.791

OpenStudy (anonymous):

oh sorry.. it is 64.79 ?

OpenStudy (anonymous):

so our solution is 65º? sorry i was in the wrong mode on my calc :O

OpenStudy (michele_laino):

that's right! make sure to set "DEG" not "RAD" on your calculator

OpenStudy (anonymous):

yay! thank you:)

OpenStudy (michele_laino):

:)